OpenStudy (anonymous):

Simplify. x^2-4x-21 all over x+3

OpenStudy (anonymous):

\[x^{2}-4x-21 = (x-7)(x+3)\]so, \[(x-7)(x+3) \div (x+3) = (x-7) \] the (x+3) cancel out leaving (x-7)

OpenStudy (anonymous):

thanks